0 of 1 pt 11 of 13 (13 complete) HW Score: 84.62%, 11 of core: 6.4.5-T Question Help Assume that females have pulse rates that are normally distributed with a mean of u 73.0 beats per minute and a stand deviation of o 12.5 beats per minute. Complete parts (a) through (c) below a. If 1 adult female is randomly selected, find the probability that her pulse rate is less than 76 beats per minute. The probability is 24 Round to four decimal places as needed.)
Answer #1

For normally distributed data, P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 73.0 beats per minute

Standard deviation = 12.5 beats per minute

a) P(X < 76 beats per minute) = P(Z < (76 - 73)/12.5)

= P(Z < 0.24)

= 0.5948

(0.5948 is the probability value corresponding to Z score of 0.24 in standard normal distribution table)